OpenStudy (anonymous):

Marcus can walk 1 mile to the library and the same time he can bicycle 2.5 miles to the sporting goods store. His speed on the bicycle is 6 mi/h faster than his walking speed. What is his speed walking?

OpenStudy (anonymous):

Lets call his walking speed 'w' and his biking speed 'b'. We know that: \[\frac{1}{w} = \frac{2.5}{b}\] WE also know that \[b=w+6\] So:\[\frac{1}{w} = \frac{2.5}{w+6}\] or: \[w+6=2.5w\] Solving for 'w' we get: \[6=1.5w\]\[w=4\] His walking speed is 4 mi/hour.